OpenStudy (moongazer):

Two sides of a right triangle measures 6 and 8, What is the third side?

OpenStudy (moongazer):

does this mean that the 2 sides measuring 6 and 8 are the legs and the third side is the hypotenuse?

OpenStudy (moongazer):

this is my interpretation of the statement

OpenStudy (jiteshmeghwal9):

Use Pythagoras theorem: - \[h^2=p^2+b^2 \implies h=\sqrt{p^2+b^2}.\] Here, p= take 6 & b= take 8 Now find "h". Tell me what do u get?

OpenStudy (moongazer):



OpenStudy (moongazer):

10

OpenStudy (jiteshmeghwal9):

Yeah , U r right dude:)

OpenStudy (moongazer):

thanks I am just clarifying it :)